### v1

# Quenched shear stresses and shear moduli in disordered elastic networks

Abstract
We consider theoretically and numerically disordered elastic networks with annealed or quenched connectivity matrices (CMs) focusing on the quenched shear stresses induced by the spontaneous symmetry breaking between different CMs. By rewriting the stress-fluctuation formalism relation for the shear modulus it is argued that theses quenched stresses give an important contribution $\mu_{1}$ to the finite shear modulus $\mu_{\mathrm{q}}=\mu_{1}+\mu_{\mathrm{a}}$ for quenched CMs. $\mu_{\mathrm{a}}$ is a simple average over all states and CMs which for equilibrium distributions of CMs is equivalent to the shear modulus of annealed networks. Interestingly, $\mu_{\mathrm{a}}$ may be strongly negative for out-of-equilibrium distributions as shown by considering a temperature quench and a scalar active two-temperature model.
Introduction.
It is generally accepted that the universe started with an extremely simple and highly symmetrical quantum state Gell-Mann (1995) which due to spontaneous symmetry-breaking fluctuations of all sorts Gell-Mann (1995); Anderson (1964); Ashcroft and Mermin (1976); Chaikin and Lubensky (1995); foo (a) say, between matter and anti-matter dominated worlds, unfolded to its present state. Broken symmetries may become permanently frozen in time, e.g., our universe should remain matter rather than anti-matter dominated, but they may also be only transiently frozen being again washed away by subsequent fluctuations after some characteristic time $t_{\star}$. Broken (permanent or transient) symmetries do not only play a role for crystalline solids Anderson (1964); Ashcroft and Mermin (1976) but are also crucial for complex matter systems without “broken continuous symmetries” but frozen local heterogeneities Götze (2009). We investigate in the present work amorphous solids formed by ideal spring networks with randomly frozen “connectivity matrices” (CMs) describing the interactions between the connected particles (“vertices”) foo (b). Due to the quenched disorder we need to be more precise concerning the averaging procedure for fluctuations George et al. (2021) being important for the determination of second derivatives of the free energy Chaikin and Lubensky (1995), especially for the shear stress relaxation modulus $\mu$ computed by means of the stress-fluctuation formalism Squire et al. (1969); Wittmer et al. (2013, 2016); George et al. (2021). We thus analyze various “moments” of the instantaneous shear stress $\hat{\tau}(c,k)$ with $k$ denoting the thermodynamic state of a CM $c$. We denote, e.g., by $\tau_{\mathrm{q}}(c)\equiv\left<\hat{\tau}(c,k)\right>_{k}$ the $k$-averaged quenched shear stress for a given $c$. The (rescaled) variance $\mu_{1}$ of these quenched shear stresses is shown to yield an important contribution to the shear modulus $\mu$.
We begin by stating the central theoretical relations assuming permanent CMs ($t_{\star}=\infty$) distributed according to an equilibrium procedure. As illustrated using a simple computational model network with annealed and quenched disorder we then generalize these relations into the time domain by allowing the reorganization of the CMs, i.e. $t_{\star}(\omega)$ becomes finite with $\omega$ being a convenient tuning parameter. This is further generalized by considering a temperature quench with an initial out-of-equilibrium distribution of CMs being slowly allowed to relax and “scalar active” networks Grosberg and Joanny (2015); Elismaili et al. (2020) with different temperatures for the vertices and the annealed CMs.
Theoretical considerations.
Let $\beta=1/T$ denote the inverse temperature (in units of Boltzmann’s constant), $V$ the system volume and $E(c,k)$ the energy of a state $k$ for a CM $c$. Following Refs. Wittmer et al. (2013, 2016); George et al. (2021) we expand $E(c,k)$ assuming a canonical transformation George et al. (2021) of the positions and momenta of the network vertices under an infinitessimal affine simple shear strain $\gamma$. The instantaneous shear stress $\hat{\tau}(c,k)$ and the instantaneous affine shear modulus $\hat{\mu}_{\mathrm{A}}(c,k)$ are defined as first and second functional derivatives of $E(c,k)/V$ with respect to this transformation Wittmer et al. (2016). While the frozen shear stresses $\tau_{\mathrm{q}}(c)\equiv\left<\hat{\tau}(c,k)\right>_{k}$ are in general finite, the ensemble of all networks has by symmetry an average shear stress
[…]
We analyze below the behavior of the (rescaled) moments
$\mu_{1}$$\equiv$ $\mu_{1}(c)\right>_{c}\mbox{ with }\mu_{1}(c)\equiv\beta V\left<\hat{\tau}(c,k)\right>_{k}^{2},$ (2)
$\mu_{2}$$\equiv$ $\mu_{2}(c)\right>_{c}\mbox{ with }\mu_{2}(c)\equiv\beta V\left<\hat{\tau}(c,k)^{2}\right>_{k},$ (3)
$\mu_{\mathrm{F}}$$\equiv$ $\mu_{\mathrm{F}}(c)\right>_{c}\mbox{ with }\mu_{\mathrm{F}}(c)\equiv\mu_{2}(c)-\mu_{1}(c),$ (4)
$\mu_{\mathrm{A}}$$\equiv$ $\mu_{\mathrm{A}}(c)\right>_{c}\mbox{ with }\mu_{\mathrm{A}}(c)\equiv\left<\hat{\mu}_{\mathrm{A}}(c,k)\right>_{k}\mbox{ and }$ (5)
$\mu$$\equiv$ $\mu(c)\right>_{c}\ \ \mbox{ with }\mu(c)\equiv\mu_{\mathrm{A}}(c)-\mu_{\mathrm{F}}(c)$ (6)
being all $\geq 0$. Due to Eq. (1) the (rescaled) variance $\mu_{1}$ is a measure of the symmetry breaking associated with the quenched shear stresses. $\mu_{\mathrm{F}}$ characterizes the typical shear-stress fluctuations for each $c$ and $\mu_{\mathrm{A}}$ the $k$- and $c$-averaged “affine shear modulus”. According to the stress-fluctuation formalism Squire et al. (1969); Wittmer et al. (2013, 2016); George et al. (2021) $\mu(c)\geq 0$ is the (thermodynamic) shear modulus of each CM $c$. Let us rewrite the $c$-averaged shear modulus $\mu$ as
$\mu=\mu_{1}+\mu_{\mathrm{a}}\mbox{ with }\mu_{\mathrm{a}}\equiv\mu_{\mathrm{A}}-\mu_{2}.$ (7)
Since $\mu_{\mathrm{F}}\geq 0$ and $\mu_{1}\geq 0$ we always have
$\mu_{\mathrm{A}}\geq\mu\geq\mu_{\mathrm{a}}.$ (8)
What is the meaning of $\mu_{\mathrm{a}}$? We denote by $l$ all combined states lumping together all $(c,k)$. Reminding also Eq. (1), the “simple average” Allen and Tildesley (2017); George et al. (2021)
$\mu_{\mathrm{a}}=\left<\hat{\mu}_{\mathrm{A}}(l)-\beta V\hat{\tau}^{2}(l)\right>_{l}$ (9)
corresponds to the formal application of the stress-fluctuation formalism disregarding the assumed symmetry breaking. To be a thermodynamically meaningful modulus not only the states $k$ of the given CMs $c$ must be at thermal equilibrium, however, but also the CMs. To see why $\mu_{\mathrm{a}}\geq 0$ for equilibrium CM-distributions let us assume that the broken symmetry becomes lifted by reversible reorganizations of the CMs for $t\gg t_{\star}$. Using Eq. (1) implies that $\mu_{\mathrm{a}}=\mu-\mu_{1}=\mu\geq 0$ is the shear modulus of annealed networks. Being a simple average $\mu_{\mathrm{a}}$ does not change if evaluated for permanently quenched CMs of same distribution albeit it differs then from the true thermodynamic shear modulus $\mu_{\mathrm{q}}=\mu_{\mathrm{a}}+\mu_{1}$ of strictly quenched CMs with finite $\mu_{1}$. ($\mu_{\mathrm{q}}$ is often used instead of $\mu$ for strictly quenched CMs.) Interestingly,
$\mu_{\mathrm{a}}\geq 0\Leftrightarrow\mu_{\mathrm{A}}\geq\mu_{2}\Leftrightarrow\mu_{\mathrm{q}}\geq\mu_{1}.$ (10)
These inequalities become equalities if the annealed system behaves for $t\gg t_{\star}$ like a liquid, i.e. $\mu_{1}=\mu=0$. For arbitrarily slowly quenched colloidal or polymer glasses kept at thermal equilibrium one thus expects Wittmer et al. (2016); George et al. (2021)
$\mu_{\mathrm{a}}=\mu_{\mathrm{A}}-\mu_{2}=0\mbox{ and }\mu_{\mathrm{q}}=\mu_{1}.$ (11)
Figure 1: Annealed elastic network with $0.1\leq R_{s}\leq 2$ and $R_{s}\right>_{s}=1$ at $T=0.1$: (a) Snapshot of a small subvolume comprising about 500 vertices. Red lines indicate compressed springs ($r<R_{s}$), blue lines extended springs. (b) Normalized distribution $p(R_{s})$.
Numerical model.
We use for illustration purposes an extremely simple two-dimensional model elastic network where $n$ vertices of mass $m$ are connected by ideal springs $s$ of energy $\frac{K}{2}(r-R_{s})^{2}$ with identical spring constants $K$ but a specific reference length $R_{s}$ for each spring. We assume $m=1$, $K=10$ and $R_{s}\right>_{s}=1$ for the first moment of the $R_{s}$-distribution. The network topology is set up using a regular hexagonal lattice of springs where each vertex is connected by identical springs of $R_{s}=1$ to its six next neighbors. At variance to the transient spring networks considered in Ref. Wittmer et al. (2016) the springs remain permanently connected to the same vertices, i.e. with this respect all CMs are quenched and identical. We use below $n=10^{4}$ vertices contained in periodic rectangular simulation boxes of linear lengths $L_{x}=100$ and $L_{y}=L_{x}\sqrt{3}/2$. The networks are sampled by means of molecular dynamics (MD) using a Langevin thermostat with a large friction constant $\zeta=1$. The standard temperature is $T=0.1$. For crystalline networks without disorder $\tau_{\mathrm{q}}(c)=0$ by symmetry for all $c$ and $T$ and, hence, $\mu_{1}=0$ and $\mu=\mu_{\mathrm{a}}$.
Annealed and quenched disorder enters by allowing $R_{s}$ to fluctuate under the additional constraint that $0.1\leq R_{s}\leq 2$. This is done by means of a Monte Carlo (MC) procedure where the reference lengths of two (randomly chosen) neighboring springs are slightly changed subject to a standard Metropolis criterion Allen and Tildesley (2017) and that both new lengths are within the indicated bounds and their sum remains constant. (This may be seen as a simple model for branched wormlike micelles with imposed number of branching points and total branch length Cates and Candau (1990).) The algorithm thus mixes MD moves for the vertices at a given CM with MC steps altering the Hamiltonian and thus the CM foo (b). Time is measured by the MD procedure where we set $\delta t=10^{-3}$ for the Verlet time increment Allen and Tildesley (2017). After each $\delta t$ we attempt for a small fraction of pairs of springs an MC hopping move to change their reference lengths. The reorganization frequency $\omega$ indicated below is proportional to this small fraction: $\omega=1$ means that a spring is attempted on average once per time unit. Results obtained for $\omega=0.01$ are shown in Fig. 1. As can be seen in panel (a), the network becomes very disordered and appears macroscopically homogeneous and isotropic. This visual impression can be corroborated using standard pair correlation functions Allen and Tildesley (2017) (not shown). Panel (b) shows the corresponding probability distribution for $R_{s}$. We thus obtain a large ensemble of $N_{\mathrm{c}}=500$ independent configurations $c$ with an equilibrium CM-distribution. Assuming ergodicity for each CM we obtain $k$-averages by means of time-averages Allen and Tildesley (2017). For instance, we estimate the $k$-average $\tau_{\mathrm{q}}(c)$ by computing the time-average $\tau_{\mathrm{q}}(c,t)$ of the instantaneous shear stresses $\hat{\tau}$ over a “sampling time” $t$. The above-introduced moments thus become apriori $t$-dependent as emphasized by an additional argument $t$ foo (c). We remind that for stationary systems the dynamical shear modulus $\mu(t)$ can be shown to be a smoothing function of the shear stress relaxation modulus $G(t)$ George et al. (2021); foo (d). Data are sampled either for strictly permanent networks ($\omega=0$) or by very gradually reorganizing the CMs at small finite $\omega$. The behavior of permanent CMs is then obtained from the large-$t$ and small-$\omega$ asymptotic limits.
Figure 2: $\mu_{\mathrm{A}}(t)$, $\mu_{2}(t)$, $\mu(t)\equiv\mu_{\mathrm{A}}(t)-\mu_{\mathrm{F}}(t)$, $\mu_{1}(t)$, $\mu_{\mathrm{F}}(t)\equiv\mu_{2}(t)-\mu_{1}(t)$ and $\mu_{\mathrm{a}}(t)\equiv\mu_{\mathrm{A}}(t)-\mu_{2}(t)$ for equilibrated and finally quenched CMs ($\omega=0$). Eq. (10) is nicely confirmed.
Equilibrium CM-distributions.
Let us first characterize the moments for switched-off reorganization MC moves ($\omega=0$). Using a half-logarithmic representation this is shown in Fig. 2 with the sampling time $t$ being the horizontal axis. As expected, all moments become rapidly $t$-independent ($t\gg 1$) as emphasized by horizontal lines. We note especially the static asymptotic limits
$\mu_{\mathrm{A}}\approx 6.2>\mu_{\mathrm{q}}\approx 4.6>\mu_{1}\approx 3.9\gg\mu_{\mathrm{a}}\approx 0.7>0.$ (12)
The inequalities Eq. (8) and Eq. (10) thus hold as expected. We also emphasize that $\mu_{\mathrm{a}}$ is very small albeit finite at variance to liquid systems, cf. Eq. (11), and the $c$-averaged shear modulus $\mu_{\mathrm{q}}=\mu_{1}+\mu_{\mathrm{a}}\approx\mu_{1}$ is thus dominated by the variance $\mu_{1}$ of the quenched stresses.
Figure 3: $\mu(t)\equiv\mu_{1}(t)+\mu_{\mathrm{a}}(t)$ for different $\omega$ as indicated in the legend. The horizontal lines indicate (top to bottom) the large-$t$ asymptotes $\mu_{\mathrm{A}}\approx 6.2$, $\mu_{\mathrm{q}}\approx 4.6$ and $\mu_{\mathrm{a}}\approx 0.7$ for $\omega=0$.
The thermodynamic static values for permanent CMs are also of relevance if the reorganization MC moves are not completely switched off. This is demonstrated for a large range of $\omega$ in Fig. 3 for $\mu(t)$. As can be seen, $\mu(t)\to\mu_{\mathrm{A}}$ for all $\omega$ for very small times (thin horizontal line) foo (c). This is expected since $\mu_{\mathrm{F}}(t)$ must vanish in this limit. For small $\omega$ and not too large $t$, $\mu(t)$ approaches the static shear modulus $\mu_{\mathrm{q}}$ of quenched networks (bold horizontal line). $\mu_{\mathrm{q}}$ becomes an intermediate upper shoulder (plateau) upon further increasing $\omega$ or $t$. Importantly, for sufficiently large $t$ and $\omega$ the dynamical shear modulus $\mu(t)$ always approaches $\mu_{\mathrm{a}}$ (dashed horizontal line), i.e. the static shear modulus of completely annealed networks. Naturally, this is best seen for our largest reorganization frequencies ($\omega\gg 0.01$). We note finally that the relaxation time $t_{\star}$, operationally defined by $\mu(t=t_{\star})/\mu_{\mathrm{a}}-1=5\%$, increases inversely with $\omega$.
Out-of-equilibrium CM-distributions.
Having confirmed Eq. (10) for equilibrium CM-distributions we show now using two examples with non-equilibrium distributions that these inequalities do not hold in general.
Figure 4: Time-dependence of observables after quenching at $t=0$ from $T=10$ to $T=0.1$ and allowing the CMs to reorganize with a small frequency $\omega=0.01$. Eq. (10) gets violated for $t\ll t_{\star}\approx 30000$ (vertical line) where $\mu_{1}(t)$, $\mu_{2}(t)$ and $-\mu_{\mathrm{a}}(t)$ are found to decay logarithmically.
Starting with annealed liquids at high temperatures amorphous solids are commonly created by a rapid quench below the glass transition temperature $T_{\mathrm{g}}$ Götze (2009). This has not only the effect that the relaxation dynamics is strongly reduced but, moreover, that glasses are generally not at thermal equilibrium at $T<T_{\mathrm{g}}$, i.e. Eq. (10) may not hold. We consider here a similar situation where our networks are first equilibrated with finite $\omega$ at $T=10$ and then suddenly at $t=0$ quenched to our standard temperature $T=0.1$, i.e. the CM-distribution is initially not at equilibrium. As shown in Fig. 4, it is then allowed to relax with $\omega=0.01$. The data points have been obtained using an iteration of subsequent tempering and measuring intervals of both $\Delta t=1000$. The indicated time for the horizontal axis is $t=\Delta t(2i-0.5)$ for each iteration step $i$. Interestingly, Eq. (10) is strongly violated for $t\ll t_{\star}\approx 30000$. For larger $t$ the equilibrium CM-distribution for $T=0.1$ is restored, cf. Eq. (12).
Figure 5: Scalar active two-temperature model with $T=0.1$ for the network vertices and $T_{\mathrm{CM}}$ for the CMs. A half-logarithmic representation of the observables as a function of the reduced temperature $x=T_{\mathrm{CM}}/T$ is used. The final production runs are performed at switched-off MC moves ($\omega=0$) sampling over a sampling time $t=10^{4}$. Eq. (10) is violated for $x\gg 1$.
As a second example we present in Fig. 5 data for two-temperature networks where the MD dynamics of the vertices is coupled to a Langevin thermostat with $T=0.1$ while a broad range of temperatures $T_{\mathrm{CM}}$ is imposed for the MC steps of the network reorganization. This model is motivated by recent work on scalar active matter with different temperatures for each particle class Grosberg and Joanny (2015); Elismaili et al. (2020). For each $x=T_{\mathrm{CM}}/T$ we anneal the CM-ensemble using a finite $\omega$ over sufficiently long times until all properties become stationary. The reorganization MC steps are then switched off for the production runs presented in Fig. 5. As expected, the data points for $x\approx 1$ are consistent with Fig. 2 and $\mu_{\mathrm{A}}(x)$, $\mu_{\mathrm{F}}(x)$ and $\mu_{\mathrm{q}}(x)$ are seen to depend very weakly on $x$. Importantly, we observe for $x\gg 1$ that the inequalities Eq. (10) are strikingly violated, just as for the short-time limit in Fig. 4. This makes sense since in both cases the CMs are distributed using a too large temperature which generates stronger quenched stresses. Moreover, Eq. (10) does not only hold for $x\approx 1$ but also for $x\ll 1$. We note finally that in the small-$x$ limit the moments $\mu_{1}(x)$ and $\mu_{2}(x)$, measuring the quenched shear stresses, slightly decrease albeit remaining finite. Hence, $-\mu_{\mathrm{a}}(x)=\mu_{2}(x)-\mu_{\mathrm{A}}(x)$ also decreases with $-\mu_{\mathrm{a}}(x)\to-3$ for $x\to 0^{+}$.
Conclusion.
We have considered in this work various contributions to the shear modulus of disordered elastic networks with annealed or quenched CMs. By rewriting the relevant relation of the stress-fluctuation formalism it was shown that the quenched shear stresses induced by the spontaneous symmetry breaking with respect to Eq. (1) yield an important contribution $\mu_{1}$ to the shear modulus $\mu_{\mathrm{q}}=\mu_{1}+\mu_{\mathrm{a}}$. The second contribution $\mu_{\mathrm{a}}$ is a simple average over all states and CMs, cf. Eq. (9), which for equilibrium CM-distributions is identical to the shear modulus of annealed networks (cf. Fig. 3). Interestingly, $\mu_{\mathrm{a}}$ may be strongly negative for out-of-equilibrium distributions as shown by considering a temperature quench (cf. Fig. 4) and a scalar active two-temperature model (cf. Fig. 5). While the shear modulus $\mu_{\mathrm{q}}$ can still be computed in these cases using the stress-fluctuation relation Eq. (7) the more special inequalities Eq. (10) do not hold.
The presented relations should also be relevant for amorphous colloidal or polymer glasses formed by a temperature quench from a liquid phase. For an extremely slow quench protocol such system may stay at equilibrium and according to Eq. (11) the static shear modulus $\mu_{\mathrm{q}}=\mu_{1}$ is set by the variance of the quenched shear stresses. More generally, however, the interacting particles may not be at equilibrium and Eq. (10) and Eq. (11) may not hold. Preliminary numerical work for colloidal glasses obtained using either rapid $T$-jumps or more gradual $T$-quenches shows indeed that $\mu_{\mathrm{a}}(T)$ becomes strongly negative below $T_{\mathrm{g}}$.
Acknowledgments.
We are indebted to J. Farago and J. Baschnagel (Strasbourg) for helpful discussions.
Data availability statement.
Data sets are available from the corresponding author on reasonable request.